OpenStudy (anonymous):

Given y = 4x^3 -12x Find the intervals where y is below x-axis Find the intervals where y is above x-axis

OpenStudy (anonymous):

Do I juz plug in random numbers and try or are there some other ways?

OpenStudy (campbell_st):

then find the roots 4x(x^2 - 3) factorised to \[4x(x - \sqrt{3})(x+\sqrt{3})\] the roots exist at \[x = -\sqrt{3} , 0. \sqrt{3}\] below the x axis \[x < -\sqrt{3}\] and \[0 < x < \sqrt{3}\]